OpenStudy (anonymous):

What is the limit of sinx/(1+cosx) as x approaches 0?

hartnn (hartnn):

directly put x=0

OpenStudy (anonymous):

That would yield 0/0

hartnn (hartnn):

1+ cos 0 =1+1 = 2

hartnn (hartnn):

or is that (1-cos x ) ??

OpenStudy (anonymous):

Right! I mistook it for a subtraction. 0/2 = 0, so the limit as x->0 is 0?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

Thank you very much!